# Standardization of Multi-Objective QUBOs

### v1

@@ Abstract
Multi-objective optimization involving Quadratic Unconstrained Binary Optimization (QUBO) problems arises in various domains. A fundamental challenge in this context is the effective balancing of multiple objectives, each potentially operating on very different scales. This imbalance introduces complications such as the selection of appropriate weights when scalarizing multiple objectives into a single objective function. In this paper, we propose a novel technique for scaling QUBO objectives that uses an exact computation of the variance of each individual QUBO objective. By scaling each objective to have unit variance, we align all objectives onto a common scale, thereby allowing for more balanced solutions to be found when scalarizing the objectives with equal weights, as well as potentially assisting in the search or choice of weights during scalarization. Finally, we demonstrate its advantages through empirical evaluations on various multi-objective optimization problems. Our results are noteworthy since manually selecting scalarization weights is cumbersome, and reliable, efficient solutions are scarce.
Index Terms:
Multi-Objective Optimization, QUBO, Standardization, Quantum Annealing.
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
Quantum computing is particularly well-suited for solving Quandratic Unconstrained Binary Optimization (QUBO) problems [1], which can be expressed as
$\min_{\boldsymbol{x}\in\mathcal{X}}\boldsymbol{x}^{\top}\boldsymbol{Q}\boldsymbol{x},$ (1)
where $\mathcal{X}=\{0,1\}^{n}$, $\boldsymbol{Q}$ is a symmetric real-valued $n\times n$ matrix, and $n$ denotes the number of binary variables, typically corresponding to qubits in quantum systems. Despite its seemingly simple quadratic form, the QUBO problem is NP-hard and serves as a unifying framework for a broad class of combinatorial optimization tasks. Notable application domains include chip design [2], finance [3, 4, 5], flight gate assignment [6], and various machine learning tasks [7, 8].
Although QUBO problems are widely applied in practice, finding an optimal solution is NP-hard [9], which makes exact optimization computationally infeasible for large instances. While exact algorithms exist [1], they typically exhibit exponential runtime in the worst case. As a result, practical approaches rely on (meta)heuristics that seek high-quality solutions efficiently, including simulated annealing [10], tabu search [11], and genetic algorithms [12].
@@
Figure 1: Plots over the distribution (b) and the objective space (a,c) of the QUBO objective functions $f(\boldsymbol{x})=\boldsymbol{x}^{\top}Q_{1}\boldsymbol{x}$ and $g(\boldsymbol{x})=\boldsymbol{x}^{\top}Q_{2}\boldsymbol{x}$. The $20\times 20$ symmetric real-valued matrices $Q_{1}$ and $Q_{2}$ are randomly generated by the qubolite package [13] with random_state seeds of 00 and $1$ respectively. Additionally $Q_{2}$ is scaled up by a factor of $10$ as well.
@@ I-A Motivation
Solution methods in MOO can be broadly categorized based on how decision-maker preferences are incorporated: no-preference, a priori, and a posteriori approaches [15, 23]. No-preference methods aim to find representative solutions without incorporating any decision-maker preferences, making them suitable when such information is unavailable or deliberately excluded. A priori methods require the decision-maker to express their preferences in advance, often through scalarization techniques that reduce the problem to a single-objective formulation [24, 25]. While effective, these methods assume that preference information is available and well-understood beforehand, which may not always be the case. In contrast, a posteriori methods first approximate or compute a diverse set of Pareto-optimal solutions, allowing the decision-maker to explore trade-offs after the optimization [26, 27, 28]. Although this reduces the need for precise prior knowledge, generating the Pareto front can be computationally intensive, especially when multiple scalarizations or high-dimensional search spaces are involved.
Given these trade-offs, we adopt a no-preference approach in this work. This choice allows us to build a simple and general foundation for demonstrating our method, without relying on strong assumptions about decision-maker preferences. Importantly, our approach can be seamlessly integrated with a priori or a posteriori techniques, which can be applied subsequently to refine the solution process based on user-defined goals or exploratory analysis.
@@ I-B Problem Setup
with $c=-\sum_{i=1}^{m}\boldsymbol{z}^{\text{ideal}}_{i}$, since the terms in the absolute values are always positive by the definition of $\boldsymbol{z}^{\text{ideal}}$. Therefore, this approach to finding a solution with no-preference is directly applicable to mQUBO as we can now just minimize the following QUBO problem:
@@ II Standardizing QUBO Objective Functions
Standardization is a commonly used method in statistics to re-scale a random variable such that it has a mean of 00 and a standard deviation of $1$. To apply it to our problem, first assume that the solution vector $\boldsymbol{x}$ can be modeled as being sampled from some probability distribution $\mathbb{P}_{\boldsymbol{X}}:\mathcal{X}\to\mathbb{R}_{[0,1]}$, where $\boldsymbol{X}$ is the random vector of binary random variables. We can then standardize the objective functions in our mQUBO problem with respect to $\mathbb{P}_{\boldsymbol{X}}$,
@@ III-A QUBO Problems
We will now lay out the mQUBO problems and methodology we will use to compare the discussed scaling methods. To start we first define 4 different QUBO problems and their respective objective functions for use in our experiments. The following max cut and subset sum problems [34] are all defined on the same set of vertices, $V(\boldsymbol{G})$:
@@
    MC$[0,1]$: A max cut problem where the weights on the graph $\boldsymbol{G}$ are sampled from the beta distribution, $\forall(i,j)\in E(\boldsymbol{G})$:
@@
    MC$\{0,1\}$: A max cut problem defined on the complete graph over the vertices of $\boldsymbol{G}$, $K_{\boldsymbol{G}}$. The weights of each edge on the complete graph are sampled from a Bernoulli distribution with equal probability of sampling 00 or $1$. However, the weight of edges that already exist in $E(\boldsymbol{G})$ are always $1$:
    $\forall(i,j)\in E(K_{\boldsymbol{G}}):\begin{cases}\boldsymbol{w}_{i,j}^{(B)}=1&(i,j)\in E(\boldsymbol{G})\\ \boldsymbol{w}_{i,j}^{(B)}\sim\mathcal{B}(0.5)&\text{otherwise}\\ \end{cases}$
@@
    MC$[1,5]$: A max cut problem defined on the complete graph over the vertices of $\boldsymbol{G}$, $K_{\boldsymbol{G}}$. The weights of each edge on the complete graph are sampled from a discrete uniform distribution from $1$ to $5$, $\mathcal{U}\{1,5\}$. However, the weight of edges that already exist in $E(\boldsymbol{G})$ are always $5$:
    $\forall(i,j)\in E(K_{\boldsymbol{G}}):\begin{cases}\boldsymbol{w}_{i,j}^{(Z)}=5&(i,j)\in E(\boldsymbol{G})\\ \boldsymbol{w}_{i,j}^{(Z)}\sim\mathcal{U}\{1,5\}&\text{otherwise}\\ \end{cases}$
@@
The max-cut problems MC$[0,1]$, MC$\{0,1\}$, and MC$[1,5]$ have objective functions with the following form [35]:
@@
These QUBO objectives can be motivated by the real-life problem of marketing a multiplayer video game and selecting a subset of influencers to be part of the marketing campaign. Here MC$[0,1]$ represents the friendship between influencers and how likely they are to get another influencer to play the game if they are part of the campaign. MC$\{0,1\}$ and MC$[1,5]$ on the other hand indicates if any two influencer share a common language and how compatible their timezones are respectively. And finally SubSum represents the budget we want to use up in the campaign.
@@ III-B Experimental setup
These scalarizations are then optimized using the digital annealer from Fraunhofer IAIS [39]. The digital annealer is given a time limit of $2$ seconds per run to solve each given scalarized problem. We then run the digital annealer $20$ times for each scalarized problem, before returning a set of solutions found over those $20$ runs. We filter this set of solutions to find the set of non-dominated solutions $\boldsymbol{S}$ found by the digital annealer. A solution $\boldsymbol{x}$ is said to dominate another solution $\boldsymbol{y}$ if every element of $\boldsymbol{x}$ is less than or equal to the corresponding element of $\boldsymbol{y}$, and at least one element in $\boldsymbol{x}$ is strictly less than the corresponding element in $\boldsymbol{y}$ [23],
@@ III-C Performance Measure: Hypervolume Indicator
Figure 3: Example of reference points (black) used for the Hypervolume calculations—that are later averaged over—in 2D objective space over problems MC$\{0,1\}$ and MC$[0,1]$.
The measure we use to compare the solutions in $\boldsymbol{S}\in\mathcal{S}$ between the different scaling methods is to compute the hypervolume of $\boldsymbol{S}$ found from optimizing the scalarizations corresponding to each method. Hypervolume essentially measures the quality and diversity of a set of non-dominated solutions by finding the size of the objective space bounded by a reference point and dominated by the objective vector of at least one solution $\boldsymbol{F}(\boldsymbol{x}),\boldsymbol{x}\in\boldsymbol{S}$ [40]. The reference point used must be dominated by all the solutions in $\boldsymbol{S}$ in objective space.
@@ III-D Results
However, for the mQUBO problem with the objectives MC$\{0,1\}$ and MC$[1,5]$, the solution found for the scalarization without any scaling performs best. That said, the hypervolume indicators for this mQUBO problem are very close to each other, with a maximum difference of $0.2962$%. We hypothesize that the reason for the similar performances is partly due to MC$\{0,1\}$ and MC$[1,5]$ being defined very similarly; they are both max-cut problems on the complete graph where edges that were already in the original graph are given the largest weight. Furthermore, from Table II, we can see that the roof dual range and standard deviation of MC$\{0,1\}$ and MC$[1,5]$ are different by just a single order of magnitude. So we believe it is a combination of both factors that led to the different approaches performing similarly.
@@ IV Conclusion
One possible avenue of future work would be to theoretically analyze the effects on the optimization landscape from scaling QUBO objectives by its standard deviation. Furthermore, it would be interesting to see how this approach of standardization can affect the integration of constraints into an mQUBO problem as well. Additionally, since we’ve only considered the variance when $\boldsymbol{x}$ is sampled from a uniform distribution, an interesting direction for future research would be to derive versions of Theorem 1 for other distributions of $\boldsymbol{X}$ and studying its impact on the pareto front, especially when subspaces of $\mathcal{X}$ become infeasible due to hard constraints.

### v2

# arXiv:2504.12419v2[cs.LG] 28 Feb 2026

@@ Abstract
Multi-objective optimization involving Quadratic Unconstrained Binary Optimization (QUBO) problems arises in various domains. A fundamental challenge in this context is the effective balancing of multiple objectives, each potentially operating on very different scales. This imbalance introduces complications such as the selection of appropriate weights to balance the different objectives. In this paper, we propose a novel technique for scaling QUBO objectives that uses an exact computation of the variance of each individual QUBO objective. By scaling each objective to have unit variance, we align all objectives onto a common scale. This allows for more balanced solutions to be found when combining these objectives directly, as well as potentially assisting in the search or choice of weights during scalarization. Finally, we demonstrate its advantages through empirical evaluations on various multi-objective optimization problems. Our results are noteworthy since manually selecting scalarization weights is cumbersome; and reliable, efficient solutions are scarce. Code for the paper can be found at https://gitlab.com/lklee/qubo-standardization.
publicationid: pubid: 10.1109/QCE65121.2025.00017 ©2025 IEEE
@@ I Introduction
A natural formulation for problems solvable by adiabatic quantum computing [1, 2] and quantum approximate optimization algorithms [3] are Quandratic Unconstrained Binary Optimization (QUBO) problems [4], expressed by:
$\min_{\boldsymbol{x}\in\mathcal{X}}\boldsymbol{x}^{\top}\boldsymbol{Q}\boldsymbol{x}=\min_{\boldsymbol{x}\in\mathcal{X}}\>\sum_{i=1}^{n}\boldsymbol{Q}_{i,i}\boldsymbol{x}_{i}+\sum_{i,j}\boldsymbol{Q}_{i,j}\boldsymbol{x}_{i}\boldsymbol{x}_{j},$ (1)
where $\mathcal{X}=\{0,1\}^{n}$, $\boldsymbol{Q}$ is a symmetric real-valued $n\times n$ matrix, and $n$ denotes the number of binary variables; typically corresponding to qubits in quantum systems. They are also equivalent to transverse-field Ising model Hamiltonians, up to the change of variables $\boldsymbol{x}_{i}=(1-\sigma^{z}_{i})/2$, where $\sigma^{z}$ is the Pauli Z operator [5]. The ground state of the resulting Hamiltonian is then equivalent to the vector $\boldsymbol{x}$ that minimizes Equation 1.
Despite its simple quadratic form, the QUBO problem is NP-hard [6] and serves as a unifying framework for a broad class of combinatorial optimization tasks. Notable application domains include chip design [7], finance [8, 9, 10], flight gate assignment [11], and various machine learning tasks [12, 13].
@@
Figure 1: Plots over the distribution (b) and the objective space (a,c) of the QUBO objective functions $f(\boldsymbol{x})=\boldsymbol{x}^{\top}Q_{1}\boldsymbol{x}$ and $g(\boldsymbol{x})=\boldsymbol{x}^{\top}Q_{2}\boldsymbol{x}$. The $20\times 20$ symmetric real-valued matrices $Q_{1}$ and $Q_{2}$ are randomly generated by the qubolite package [14] with random_state seeds of 0 and 11 respectively. Additionally $Q_{2}$ is scaled up by a factor of $10$ as well.
@@ I-A Motivation
Solution methods in MOO can be broadly categorized based on how decision-maker preferences are incorporated: no-preference, a priori, and a posteriori approaches [16, 24]. No-preference methods aim to find representative solutions without incorporating any decision-maker preferences, making them suitable when such information is unavailable or deliberately excluded. A priori methods require the decision-maker to express their preferences in advance, often through scalarization techniques that reduce the problem to a single-objective formulation [25, 26]. While effective, these methods assume that preference information is available and well-understood beforehand, which may not always be the case. In contrast, a posteriori methods first approximate or compute a diverse set of Pareto-optimal solutions, allowing the decision-maker to explore trade-offs after the optimization [27, 28, 29]. Although this reduces the need for precise prior knowledge, generating the Pareto front can be computationally intensive.
In this work, we adopt a no-preference approach. This choice allows us to build a simple and general foundation for demonstrating our method, without relying on strong assumptions about decision-maker preferences. Importantly, our approach can be seamlessly integrated with a priori or a posteriori techniques, which can be applied subsequently to refine the solution process based on user-defined goals or exploratory analysis.
@@ I-B Problem Setup
with $c=\sum_{i=1}^{m}\boldsymbol{z}^{\text{ideal}}_{i}$, since the terms in the absolute values are always positive by the definition of $\boldsymbol{z}^{\text{ideal}}$. Therefore, this approach to finding a solution with no-preference is directly applicable to mQUBO as we can now just minimize the following QUBO problem:
@@ II Standardizing QUBO Objective Functions
Standardization is a commonly used method in statistics to re-scale a random variable such that it has a mean of 0 and a standard deviation of $1$. To apply it to our problem, first assume that the solution vector $\boldsymbol{x}$ can be modeled as being sampled from some probability distribution $\mathbb{P}_{\boldsymbol{X}}:\mathcal{X}\to\mathbb{R}_{[0,1]}$, where $\boldsymbol{X}$ is the random vector of binary random variables. We can then standardize the objective functions in our mQUBO problem with respect to $\mathbb{P}_{\boldsymbol{X}}$,
@@ III-A QUBO Problems
We will now lay out the mQUBO problems and methodology we will use to compare the discussed scaling methods. To start we first define 4 different QUBO problems and their respective objective functions for use in our experiments. The following max cut and subset sum problems [35] are all defined on the vertices $V(\boldsymbol{G})$ of a random Barabasi-Albert graph $G$. Each QUBO problem utilizes the connectivity of the initial graph $G$ in its construction.
@@
    MC[0,1][0,1]: A max cut problem where the weights on the graph $\boldsymbol{G}$ are sampled from the beta distribution, $\forall(i,j)\in E(\boldsymbol{G})$:
@@
[…]
@@
    MC[1,5][1,5]: A max cut problem defined on the complete graph over the vertices of $\boldsymbol{G}$, $K_{\boldsymbol{G}}$. The weights of each edge on the complete graph are sampled from a discrete uniform distribution from $1$ to $5$, $\mathcal{U}\{1,5\}$. However, the weight of edges that already exist in $E(\boldsymbol{G})$ are always $5$:
[…]
@@
The max-cut problems $MC{0,1}\{0,1\}$, and MC[1,5][1,5] have objective functions with the following form [36]:
@@
[…]
@@ III-B Experimental setup
These scalarizations are then optimized using the digital annealer from Fraunhofer IAIS which implemets an evolutionary algorithm for solving QUBO problems on an FPGA [40]. The digital annealer is given a time limit of $2$ seconds per run to solve each given scalarized problem. We then run the digital annealer $20$ times for each scalarized problem, before returning a set of solutions found over those $20$ runs.
We filter this set of solutions to find the set of non-dominated solutions $\boldsymbol{S}$ found by the digital annealer. A solution $\boldsymbol{x}$ is said to dominate another solution $\boldsymbol{y}$ if every element of $\boldsymbol{x}$ is less than or equal to the corresponding element of $\boldsymbol{y}$, and at least one element in $\boldsymbol{x}$ is strictly less than the corresponding element in $\boldsymbol{y}$ [24],
@@ III-C Performance Measure: Hypervolume Indicator
[…]
The measure we use to compare the solutions in $\boldsymbol{S}\in\mathcal{S}$ between the different scaling methods is to compute the hypervolume of $\boldsymbol{S}$ found from optimizing the scalarizations corresponding to each method. Hypervolume is a standard approach for evaluating solution sets to MOO problems and essentially measures the quality and diversity of a set of non-dominated solutions. It does so by finding the size of the objective space dominated by the objective vector of at least one solution $\boldsymbol{F}(\boldsymbol{x}),\boldsymbol{x}\in\boldsymbol{S}$ and bounded by a chosen reference point [41]. The reference point used must be dominated by all the solutions in $\boldsymbol{S}$ in objective space. Therefore, the larger the hypervolume indicator is, the greater the region the solution set $\boldsymbol{S}$ dominates in objective space; implying a solution set of higher quality.
@@ III-D Results
[…]